# Vistoria: A Multimodal System to Support Fictional Story Writing through Instrumental Text–Image Co-Editing

### v2

@@ 6.1.4. Structural Mapping across Multiple Dimensions to Trace the Story
Figure 12. Images act as cognitive scaffolds, helping writers describe unfamiliar actions or contexts more concretely. P5 constructed her final writing by referring to both images and adopted and edit related text.
6.2. Multimodal Approaches Creation to Inspire Creativity
To answer RQ2, we found that image-text co-editing facilitated creativity in fictional story writing through three key cognitive processes: (1) divergent exploration, where participants spatially arranged multimodal components and pursued multiple narrative directions; (2) serendipitous discovery, where unexpected LLM outputs sparked new creative possibilities; and (3) multimodal cues activating memory and cultural connections. Together, these processes enhanced participants’ sense of authorship by positioning the system as a collaborative partner that expanded creative possibilities rather than replacing human imagination.
@@ 6.2. Multimodal Approaches Creation to Inspire Creativity
6.2.1. Bottom-up Creation to Support Divergent Exploration
As shown in Figure 13 and Table 3, participants demonstrated significantly greater breadth and depth of exploration with Vistoria compared to the GPT baseline. They pursued a wider range of narrative directions ($M_{\text{System}}=6.92,SD=2.81$ vs. $M_{\text{Baseline}}=1.42,SD=1.08$) while also producing more variations within single paths (branches: $M_{\text{System}}=3.00,SD=1.35$ vs. $M_{\text{Baseline}}=1.92,SD=0.67$). Although the average depth per path was slightly lower for Vistoria ($M_{\text{System}}=1.70,SD=1.18$ vs. $M_{\text{Baseline}}=2.00,SD=1.22$), this reflected participants’ substantially broader exploration overall; within individual directions, they often engaged in deeper iterations, as evident in Figure 13. Such enhanced exploration transformed their creative process from passively consuming LLM outputs to actively curating multimodal components.
The Vistoria canvas functioned as an exploratory space where participants pursued multiple storylines in parallel without linear constraints. This freeform environment encouraged spatial arrangement, clustering, and repositioning of ideas through manipulation and recombination of alternative story–image cards to explore divergent scenarios (P7, P4, P5, P8). Rather than committing immediately to single narratives, participants typically generated multiple alternatives in early phases, positioning the system as an expressive medium rather than merely a text generator. This exploratory approach helped participants avoid early fixation and sustained creative engagement. As P8 observed: “GPT workflow is more streamlined… top-down. Using the visual app feels more bottom-up. You are open to possibilities, so there’s not a finite result and more possibilities being explored.” Ultimately, this shift highlights how multimodal, spatially organized exploration fosters divergent thinking, sustains creative momentum, and supports a more generative form of narrative authorship.
6.2.2. Serendipity from LLM Randomness Enriches Narrative Development
Another recurring theme was that the inherent randomness of LLM-generated image-text output often introduced serendipitous elements into participants’ creative process. Rather than being perceived as noise, these unexpected details were frequently embraced as inspiration, enriching story development and occasionally opening entirely new narrative directions.
For example, when P1 generated an image-text pair of the mysterious box, the system unexpectedly inserted the logo “LEH.” on the box and in the text description. Intrigued, P1 subsequently asked the system to visualize who or what “LEH” could be, gradually evolving this accidental motif into a recurring symbol within their final story (Figure 11). Nearly all participants described how surprises in LLM-generated images or texts gave them fresh possibilities to explore. P12 captured this dynamic vividly: “The system gave me a riddle at an early stage, I didn’t understand it at first, but as I kept working, I slowly decoded it, and this process of unraveling became part of my story. It helped me construct something unique that I wouldn’t have created otherwise.” (Figure 14) Together, these examples show how the interplay of multimodal generation and LLM’s inherent unpredictability fueled creativity. Randomness supported divergence by surfacing novel ideas, while participants retained agency in selectively integrating these elements, leading to unique, personally authored narratives.
Figure 14. In the creation process of P12, unexpected LLM randomness served as serendipitous prompts: a generated riddle was decoded and transformed into a key-and-note motif guiding Claire toward the lighthouse, illustrating how stochastic details became generative cues that enriched development and redirected the plot.
6.2.3. Unlocking Creativity Through Memory and Cultural Resonance
Participants reported that multimodal cues routinely triggered personal memories and cultural references, enriching their stories. Visual props, e.g., a medieval mask, readily invoked film/literary associations; as P4 noted, “Seeing the mask immediately brought Batman/Phantom to mind… it made choosing a style much easier.” (Figure 15) Crucially, image–text pairs co-triggered lived experiences. In P7’s case, a fountain scene image together with the phrase “mossy stone” jointly evoked a vivid memory, which she then translated into a concrete narrative element, the envisioned fall that became her story’s ending(Figure 15). These multimodal traces turned associative recall into usable story material, weaving authenticity and imaginative depth into the narrative.
6.2.4. Strengthened Authorship and Agency
Enhanced exploration with Vistoria strengthened participants’ sense of authorship. While GPT outputs in the baseline condition were consistently described as “surface-level” (P2, P5, P7) and felt like “someone else’s work” (P3), almost all participants reported markedly stronger authorship with Vistoria, enabling deeper exploration and richer plots (P1, P2, P3, P5, P12, P10). P5 articulated this shift:
“When using Vistoria, every idea originated from my own imagination, and the final story was formed by combining these different self-generated ideas. This gave me a strong sense that the story was truly my own creation. I think of the system less as a generator and more as a companion sketchbook. While you explore different directions, you always remain in charge of the final piece.”
P9 reinforced this sentiment: “This tool is more ‘me’… I control characters and plots; ChatGPT may not become my story.” Quantitative results corroborated these perceptions: participants rated Vistoria as providing stronger collaborative experiences (CSI: $M_{\text{Vistoria}}=5.50$ vs. $M_{\text{Baseline}}=4.58$, p = .0418) (Table 2). Participants framed Vistoria as a supportive co-pilot rather than a replacement; as P10 noted: “I consider this system as a copilot that does not help you with everything, and this system that you have to create by yourself.” Together, these findings mark a transition from passively adopting others’ ideas to actively curating one’s own generative outputs, reinforcing creative agency and ownership.
Figure 15. Evolving cards leave visual traces that cue personal memories and cultural references (e.g., a mask → Batman/Phantom; a “mossy stone” → place memory), which writers externalize into concrete scenes. These multimodal traces act as mnemonic scaffolds, helping retrieve earlier thinking and weave lived experience into the story world—transforming associative recall into narrative material.
6.3. Tradeoffs of Multimodal Interaction for Creative Writing
While image–text co-editing enhanced authorship, coherence, and expressiveness, it also increased workload. NASA–TLX scores showed significantly higher mental demand ($M_{\text{Vistoria{}}}=5.16$ vs. $M_{\text{baseline}}=3.17$, p=.0000) and physical demand ($M_{\text{Vistoria{}}}=4.67$ vs. $M_{\text{baseline}}=2.08$, p=.0002), plus moderately higher frustration ($M_{\text{Vistoria{}}}=2.75$ vs. $M_{\text{baseline}}=1.75$, p=.0204) (Table 2). These increases reflect the effort of coordinating across modalities and actively curating outputs relative to the GPT baseline. Part of the mental load may have stemmed from first-time use—learning new image/text operations and switching between modalities (P1, P2, P5, P9, P10): “The biggest burden is switching between tools to sketch or type; it takes time to learn and adapt, even though the functions are useful.” Additional difficulty arose from unfamiliarity with the canvas interface (P3, P10).
6.3.1. Higher burden of authorship.
Almost all participants valued the ability to maintain control of the story, and the creation process gave them a stronger sense of ownership. Yet, this empowerment came with a cost: several participants noted that they had to actively develop details within their own text, especially at the early stages, which could feel “a little bit frustrating” (P9). Unlike GPT, which could quickly produce long passages or propose questions to guide brainstorming (P10), the system required users to supply and elaborate on their own ideas before meaningful generation occurred, which may lead to a higher mental workload.
This shift demanded more cognitive and physical effort: even though sketching and annotation helped externalize mental imagery, participants noted that it felt more demanding than simply receiving GPT’s ready-made text (P1, P2, P11). Thus, stronger authorship agency came at the cost of higher workload.
6.3.2. Validating ideas rather than generating them.
Some participants noted that the system’s strengths lay in validating or expanding concrete mental images rather than generating abstract directions(P1, P11). As P1 explained, “when I have a vague impression in my mind, I tend to generate some image-text pairs. But sometimes, once the visual appears, it fixes my imagination in a certain way in my mind, and I can no longer imagine other possibilities. In contrast, only plain text can inspire limitless imagination.” This reveals a tension: images act as concrete anchors that aid detailed development, yet their representational specificity can induce fixation by prematurely crystallizing fuzzy concepts and narrowing exploration.
Similarly, P10 noted that the baseline GPT condition was superior at breaking down initial story points and offering prompt-like guidance, whereas the multimodal Vistoria system primarily served to elaborate or diverge from existing visions. This made the multimodal approach particularly valuable for writers with partially formed concepts, but less helpful during the most open-ended phases of ideation when abstract exploration is more important than visual specificity.
6.3.3. Cognitive offloading and efficiency despite higher load.
Despite a higher mental workload, participants repeatedly emphasized that multimodality enabled more efficient allocation of attention. From participants’ view, the image-text pairs externalized fleeting ideas, preserved spatial and sensory detail, and reduced information loss when translating imagination into tangible artifacts (P4, P7). As P7 noted, “By highlighting and collaging, I externalized formed ideas into image–text pairs, clearing mental space to pre-plan the next line and concentrate on the next plot beat.”
Thus, while Vistoria required more effort and decision-making, it also freed participants from lower-level memory maintenance and allowed them to focus on higher-order creative synthesis. With greater familiarity, the efficiency gains from this offloading may outweigh the initial overhead.
7. Discussion
7.1. Cross-modal LLM-Mediated Instrumental Interaction
We extend the notion of Instrumental Interaction into an LLM-mediated, cross-modal paradigm where generated artifacts repeatedly re-enter the workflow as manipulable outputs. In classic formulations, instruments mediate between users and domain objects, maintaining a separation between “tool” and “content” (Beaudouin-Lafon, 2000). In generative contexts, this dichotomy breaks down: each card (image–text pair) is simultaneously an outcome and a renewed object of operation, enabling recursive, non-linear composition (Riche et al., 2025).
This reframing also reshapes reification. Vague creative intents expressed through text, sketches, or images materialize across modalities into manipulable, cross-linked artifacts that remain editable and reusable. Reification thus spans modalities rather than existing in a single channel, supporting direct manipulation and reflection on both structure and semantics of evolving narratives (Riche et al., 2025; Masson et al., 2024b).
A key mechanism enabling this circulation is structural mapping (Gentner, 1983; Norman, 2013; Masson et al., 2024b). Participants made sense of their evolving materials not only by inspecting the surface properties of each medium but also by exploiting the correspondences in shared properties, concepts, or identities established between them. Refining a visual scene, for example, carried implications for narrative structure, tone, or pacing in text, while textual edits translated into compositional adjustments in images. In this way, mappings transformed what is usually a cognitive act of alignment into an operational resource: they stabilized coherence while keeping alternatives visible and accessible. Instrumental interaction and structural mapping thus operate as mutually reinforcing processes. Instruments generate the artifacts that mappings align, while mappings scaffold the interpretability and reuse of instruments.
This dynamic carries several design implications for future systems. They should not only offer isolated multimodal tools but cultivate instrumental ecosystems in which mappings are first-class and inspectable, and where families of instruments carry coherent semantics across domains (Riche et al., 2025). Furthermore, instead of overwhelming users with immediate complexity, operations can be gradually disclosed, allowing mappings to surface progressively and adapt to context (Shneiderman and Maes, 1997; Allen et al., 1999). Provenance views that expose the cascade of edits across modalities could support both divergence and recombination, helping individuals and teams understand how narratives branch and converge (Masson et al., 2024b).
Ultimately, the LLM can be positioned less as a generator of content and more as a mediator of instruments that align structures, propagate semantics, and synthesize new operations from prior materials. This shift reimagines authorship as an orchestration of evolving, cross-modal instruments, opening a design space where exploration and coherence can be sustained simultaneously (Sawyer, 2017; Masson et al., 2024b).
7.2. Orchestration and Curatorial Authorship to Form Creativity in Cross-Modal Creation
Traditional creativity theories emphasize individual expression through language and pre-formed intent (Smith, 2020; Donnelly, 2011), positioning the creator as the primary originator of content. Our findings challenge this view by demonstrating that creativity in Vistoria manifested as orchestration. Participants conducted and harmonized streams of multimodal LLM outputs through selection, combination, and reframing. Creative goals emerged through interaction with generated materials rather than preceding them, with meaning composed across representational boundaries. This suggests creativity in AI-augmented contexts should be theorized as the strategic coordination of evolving possibilities, repositioning authorship from origination to curation and synthesis within a collaborative ecosystem.
While existing multimodal creativity research often focuses on seamless integration between modes (Yen et al., 2025; Rosenberg et al., 2024; Shi et al., 2025), our study reveals that productive tensions between text and image became central to the creative process. Participants actively negotiated the trade-offs between imagistic specificity and narrative openness, treating these tensions not as problems to solve but as resources for exploration. The ongoing adjudication process—accept, adapt, defer, or discard–became the primary engine of creative coherence. This extends theories of creative constraint by showing how representational tensions across modalities can drive rather than hinder creative development.
Furthermore, our findings contribute to debates about human agency in LLM-assisted creation by demonstrating that curatorial work enhanced rather than diminished participants’ sense of ownership. Authorship was experienced as steering a dialogue with the model, including deciding what counts, where elements belong, and how disparate fragments accrue into a unified voice. This sense of agency grew precisely from having meaningful choices to negotiate among viable alternatives (Resnick et al., 2005). This suggests a new model of distributed authorship where human control emerges through deliberate curation within a space of LLM-generated possibilities.
As for design implications, our findings suggest that future system design should prioritize support for orchestration over seamless, one-shot generation. Instead of producing singular outputs, tools could by default surface multiple cross-modal alternatives, accompanied by lightweight mechanisms that enable users to accept, adapt, defer, or discard options as part of an ongoing multimodal curatorial process (Resnick et al., 2005). Moreover, systems should not only present outputs but also make the inherent tensions between modalities visible and adjustable. This can be done, for example, through interactive controls such as ambiguity sliders that allow users to navigate between imagistic specificity and narrative openness.
Beyond simple interface controls, future systems could also provide deeper support for curatorial authorship. For example, they might keep track of where each fragment or idea came from across different iterations, or offer reusable “curation moves” that let users quickly repeat and adapt common ways of combining or editing materials (Shneiderman, 2007). In this way, design emphasis shifts away from achieving frictionless integration and toward helping users actively steer evolving creative possibilities, foregrounding authorship in AI-mediated contexts as a matter of curation, synthesis, and negotiated choice (Ding, 2024).
7.3. Limitation and Future Work
Task Scope and Short-Term Focus. The 300–-500 word story task, while manageable for controlled evaluation, does not reflect the demands of long-form fictional writing, where authors build sustained voices, complex arcs, and intricate structures. We did not have an opportunity to study Vistoria’s suitability for extended projects, iterative revisions, or complex narratives, leaving open questions on the consistency in long works, scalability with larger volumes, or risks of over-reliance on LLM over the time of use. Moreover, evaluation relied primarily on self-reports of creativity and user experience; we did not include objective measures of story quality, originality, or literary merit.
In the future, we plan on conducting field studies to deploy this system with writers of varying expertise levels in their authentic creative contexts, observing how they integrate the tool into real writing projects over extended periods. Such longitudinal research would assess the ecological validity of Vistoria and provide insights into how writers adapt Vistoria for planning and translating across longer creative cycles. We anticipate that writers might spontaneously capture inspirational moments from daily life, potentially increasing their reliance on clustering functionality as they generate more dispersed content fragments that require organization. These naturalistic studies would provide crucial insights into the tool’s role in sustained creative practice, revealing usage patterns, adaptation strategies, and long-term impacts on writers’ creative processes that controlled laboratory settings cannot capture.
Participant Sample. Our study involved only 12 participants, while this is typical for similar lab usability studies, a larger group could provide stronger statistical power, reveal more varied interaction patterns, and allow comparisons across subgroups. Furthermore, the participant pool has limited cultural and age diversity, which could have narrowed the range of narrative traditions, writing styles, and storytelling approaches represented. Future research should address these limitations by recruiting a larger and more diverse set of participants, including professional authors, older and younger writers, and individuals from varied cultural backgrounds, to more fully evaluate the system’s applicability and generalizability.
8. Conclusion
This paper introduced Vistoria, a multimodal text–image co-editing system designed to support the cyclical “planning–translation” process in fictional writing. Grounded in Wizard-of-Oz co-design study and theories of Instrumental Interaction and Structural Mapping, the system integrates four key capabilities: (1) representing sketches, texts, and references as manipulable card units; (2) enabling cross-modal alignment so that edits in one modality propagate to the other; (3) providing polymorphic operations that apply consistently across text and image; and (4) supporting organization and reuse through highlighting and clustering panels. A controlled study with 12 creative writers demonstrated that Vistoria enhanced writers’ expressiveness, immersion, and sense of collaboration, while broadening exploration of narrative branches and reinforcing authorship and control. These findings suggest that Vistoria transforms the writing process into a dialogic interaction with multimodal instruments, enabling authors to negotiate meaning and control across modalities.
Appendix A Appendix
A.1. Writing Topics Used in the User Study

### v3

@@ 6.2.2. Bottom-up Creation to Support Divergent Exploration
Figure 10. Images act as cognitive scaffolds, helping participants describe unfamiliar actions or contexts more concretely. P5 constructed her final writing by referring to both images and adopting and editing related text.
The figure shows an example canvas during early exploration, where multiple image–text cards, sketches, and annotations are distributed spatially. Different narrative directions are pursued in parallel, illustrating non-linear exploration before committing to a single storyline.
6.2.3. Leveraging Image-Text Alignment for Rich Descriptions and Story Progression
When text and images appear together, the two modalities reinforce one another, enabling unfamiliar or imagined elements to be both visualized and verbalized. This cross-modal grounding supports the production of more detailed and enriched descriptions.
Visual references, in particular, helped participants imagine actions or settings beyond their lived experience. As P5 noted, “Looking at an image makes it easier to kind of, like, describe what something is supposed to look like or feel like… so when you describe it to [readers], and they can picture it for themselves in their head, that is what makes the story more interesting.” From observation, P5 closely described the scene in her imagination with the cliff, the necklace, and the cat in the image in her writing and directly adopted some text generated, integrating them into her final story (Figure 10).
Using the generated text and image also helped produce more concrete, detailed narratives. For example, P1 used visual cues from an image showing Claire touching a letter and adopted the descriptions in the text, such as “Claire steadies the box on her lap” and created the narrative of “She runs her fingers over the letters, heartbeats echoing in her ears.” to form his final story. Here, the text description with visualization allow participant to capture more dynamic, sensory-rich narrative moments.
6.2.4. Preserving Sense of Agency and Ownership.
Unlike the baseline, where participants often felt like they were editors of GPT-generated text, Vistoria supported exploratory editing while preserving a sense of agency and ownership. Participants felt that Vistoria “having more sense of mastery over the content,” in contrast to the baseline (P4, P6, P9). P7 felt that when using Vistoria, she was actively cutting, combining, filtering, and directing the story’s trajectory, whereas with GPT, she was mostly receiving and editing what the model produced. P9 explicitly emphasized a heightened sense of agency, noting: “This tool is more ‘me’… I control characters and plots.”
Dissatisfied with the baseline condition, in which GPT produced most of the content and left participants primarily in the role of adopting it (P4, P6, P7), P10 also characterized Vistoria as a supportive co-pilot rather than a substitute for their own work. Quantitative results are consistent with these perceptions: participants rated Vistoria as providing a stronger collaborative experience than the baseline (CSI: $M_{\text{Vistoria}}=5.50$ vs. $M_{\text{Baseline}}=4.58$, p = .0418; Table 1).
This preserved sense of agency also led to the preserved sense of ownership. GPT outputs in the baseline condition were repeatedly described as “surface–level” (P2, P5, P7) and as “someone else’s work” (P3), whereas most participants reported a stronger sense of ownership with Vistoria (P1, P2, P3, P4, P5, P6, P8, P9, P10, P11). As P5 explained: “When using Vistoria, every idea originated from my own imagination, and the final story was formed by manipulating and combining these different self-generated ideas. This gave me a strong sense that the story was truly my own creation.”
Together, these findings suggest a shift from passively adopting model suggestions to actively creating and curating one’s own generative outputs.
7. Discussion
7.1. Multimodal Instrumental Interaction
Instrumental Interaction (Beaudouin-Lafon, 2000) conceptualizes instruments as mediators that translate writers’ actions into operations on domain objects. We operationalize this principle by reifying a set of multimodal instrumental operations (Collage, Lasso, Filter, and Perspective Shift) that simultaneously act upon both text and image narrative materials. Rather than treating images and text as separate interface elements, these instruments serve as unified interactional units that enable writers to zoom between narrative scales, reorganize multimodal story fragments, and explore divergent directions within a shared representational space to edit image and text simultaneously.
From the perspective of Designing Interaction, not Interfaces (Beaudouin-Lafon, 2004), our design moves away from adding more interface widgets and focuses instead on shaping the quality of writers’ ongoing activity. Beaudouin-Lafon argues that transformative interfaces must shift attention from surface-level user-interface (UI) components toward the underlying interactional structures that support creative work (Beaudouin-Lafon, 2004). Following this perspective and Dual Coding Theory’s suggestions for verbal and nonverbal perception alignment (Clark and Paivio, 1991), Vistoria’s design prioritizes fluid transitions between modalities, persistent manipulable artifacts, and an iterative loop in which narrative ideas and multimodal materials co-evolve. This interaction-centered framing explains why writers perceived the system as increasing expressiveness.
Although participants generally found the instruments effective, they also reported a substantial learning curve. For novice system writers, the system imposed considerable mental and physical workload. These insights suggest future design opportunities to lower workload and adapt to writers’ evolving needs. In the early stages of system use, writers could express their intentions through natural language, allowing the system to suggest the most appropriate functions on their behalf (Bourgault et al., 2025). As writers become more familiar with the system, they can choose functions by themselves and even define customized functions that better fit their evolving needs. This approach aligns with the emphasis on supporting diverse, situated practices rather than enforcing a fixed interface vocabulary (Beaudouin-Lafon, 2004). Together, these directions point toward a design space where multimodal creativity systems integrate explicit instruments with adaptive, activity-centered interaction models to better support real-world writing workflows.
7.2. Designing Dynamic Mixed-Initiative Multimodal Workflows
In the traditional turn-based GPT workflow, writers often occupy a relatively passive or evaluative role: they receive model output and act primarily as examiners who check, accept, or correct the result (Kwan et al., 2024). In Vistoria, writers actively manipulate elements on the surface, decide which multimodal materials to combine, and select which operations to apply. From the perspective of the participants, Vistoria is not experienced as a detached “answer engine,” but as a co-pilot collaborator. Writers perceive themselves as those who decide how to create, what to keep, and which tools to invoke, preserving the sense of agency and ownership.
However, this also brings cognitive effort. Precisely because the workflow is instrument and manipulation-driven, it demands that writers have a clearer sense of what they need, or at least which direction they wish to explore. When writers do not yet know what they want, the system requires them to specify intentions and choose operations that tend to lead to higher cognitive load. This reflects an inherent tension: high functional agency (agency is structurally distributed in the system) requires clear intentionality, but early-stage ideation is often characterized by ambiguity and exploration (Rafner et al., 2025). In contrast, a text-only GPT chat enables the rapid generation of a large amount of content, allowing for subsequent refinement through iterative prompting and selection to achieve a specific focus. This suggests that systems should adapt their agency distribution across creative phases as agency and ownership are not static end-states but fluctuate across the co-creation trajectory (Rafner et al., 2025). In early divergent exploration, lower functional agency may paradoxically support higher felt agency (e.g., through control, decision-making, or ownership) by reducing decision fatigue; in later convergent refinement, higher functional agency (granular control) aligns with writers’ desire to shape details (Rafner et al., 2025; Fu et al., 2024).
On the other hand, ownership emerges from users’ active control over outputs and their ability to modify content (Draxler et al., 2024a). Critically, this control includes not only refinement flexibility but also rejection. Our results show that participants sometimes think image output “fixed imagination in a certain way,” which may threaten to constrain their creative vision. Previous work indicates that when AI outputs deviate from expectations, users experience diminished control and ownership uncertainty (Draxler et al., 2024a), and encouraging system design to prompt them to deploy adaptive strategies to restore agency. In that case, future designs could operationalize this through a function that, when invoked, asks “What specifically do you want to avoid?” and generates alternatives that explicitly diverge from the rejected output. This would transform counter-inspiration (Rafner et al., 2025) from an ad-hoc user strategy into operations that reframe dissatisfaction from a null action into a mechanism for articulating and preserving authorial boundaries and ownership.
7.3. Limitation and Future Work
Task Scope and Short-Term Focus. The 300-to-500-word story task, while manageable for controlled evaluation, does not reflect the demands of long-form fictional writing, where authors build sustained voices, complex arcs, and intricate structures (Chen et al., 2025). We also did not have an opportunity to study Vistoria’s suitability for extended projects, iterative revisions, or complex narratives, leaving open questions on the consistency in long works, scalability with larger volumes, or risks of over-reliance on LLM over time of use. Moreover, evaluation relied primarily on self-reports of creativity and user experience; we did not include objective measures of story quality, originality, or literary merit.
In the future, we plan to conduct field studies that deploy this system with writers of varying expertise levels in their authentic creative contexts, observing how they integrate the tool into real writing projects over extended periods. Such longitudinal research would assess the ecological validity of Vistoria and provide insights into how writers adapt Vistoria for planning and translating across longer creative cycles (Long et al., 2025). We anticipate that writers might spontaneously capture inspirational moments from daily life, potentially increasing their reliance on clustering functionality as they generate more dispersed content fragments that require organization. These naturalistic studies would provide crucial insights into the tool’s role in sustained creative practice, revealing usage patterns, adaptation strategies, and long-term impacts on writers’ creative processes that controlled laboratory settings cannot capture.
Participant Sample. Our study involved only 12 participants, while this is typical for similar lab usability studies, a larger group could provide stronger statistical power, reveal more varied interaction patterns, and allow comparisons across subgroups. Furthermore, the group of participants has limited cultural and age diversity, which could have narrowed the range of narrative traditions, writing styles, and storytelling approaches represented. Future research should address these limitations by recruiting a larger and more diverse set of participants, including writers of various ages and individuals from diverse cultural backgrounds, to more fully evaluate the applicability and generalizability of the system.
Construct Validity and Measurement Limitations. Although we discuss constructs such as creativity, sense of ownership, and agency, these observations arise primarily from qualitative reports. Our study does not include construct-grounded measurements or comparative baselines for these phenomena. Accordingly, the interpretations should be viewed as exploratory insights rather than empirically validated effects. Future work will incorporate construct-aligned measures and validated scales such as the Mixed-Initiative CSI (Lawton et al., 2023), or using frameworks situated within human–AI co-creativity like COFI (Rezwana and Maher, 2023) and CCDF (Davis et al., 2025) to quantify co-creativity interaction dynamics or better derive Human-AI co-creativity insights (Rezwana and Maher, 2025).
Furthermore, while we assessed participants’ AI usage frequency (e.g., daily ChatGPT usage), we did not measure their collaborative AI literacy or metacognitive capabilities when working with AI systems. Future research should incorporate validated scales  (Sidra and Mason, 2025) to investigate how collaborative AI literacy influences the effectiveness of multimodal co-editing tools.
Multimodality Scope. Our work focuses on multimodal support through text and images, but does not include other modalities. Prior research in creative writing suggests that audio can also serve as a useful medium (Singh et al., 2023), especially through nonverbal sounds and ambient effects that help shape mood and atmosphere. In future work, we plan to explore the addition of audio cues to the writing process. Such sound elements may support writers in building a stronger vibe, enhancing scene-setting, and offering an additional channel for creative inspiration (Rezwana and Maher, 2023).
8. Conclusion
This paper presents Vistoria, a multimodal image–text co-editing system that supports fictional story writing by tightly integrating image and text representations. Grounded in the WoZ co-design study, Vistoria introduces a unified set of instrumental operations (Lasso, Collage, Perspective Shift, and Filter) that reify writers’ intentions and enable synchronized manipulation across modalities. Through a controlled user study, we demonstrate that multimodal co-editing enhances expressiveness, immersion, and exploratory ideation. Although this multimodal workflow increases cognitive demand, participants reported preserved senses of agency and ownership, treating the system as a creative partner rather than a generative tool. We hope Vistoria highlights the opportunities for designing future writing systems that embrace multimodality as a core mechanism for ideation and narrative development.
Acknowledgments
This work is supported by City University of Hong Kong Teaching Development Grant (6000901), City University of Hong Kong Booster Fund (7030021), and Hong Kong Research Grants Council Theme-based Research Scheme (T45-205/21-N). We sincerely thank all anonymous reviewers for their insightful comments and constructive feedback that helped improve this paper.
Appendix A Appendix
A.1. User Behavior in the WoZ co-design study
Figure 11. Results of user strategies for manipulating multimodal elements in the WoZ co-design study.
Results of user strategies for manipulating multimodal elements in the WoZ co-design study.
A.2. Behavioral interaction data gathered from participants in the user study
@@
The figure compares story development outcomes between the baseline text-only condition and Vistoria. It visualizes differences in narrative branching and structural diversity, showing how Vistoria supports parallel exploration and alternative story trajectories.
We examined the sequences of function use and compared exploration patterns across the baseline and our condition. To characterize participants’ divergent–convergent behaviors during the creative process, for each task, we reconstructed exploration structures by defining Directions as top-level trajectories toward a goal, Branches as the diversity of possibilities generated within a direction, and Depth as the mean number of iterative steps within each branch to compare the exploration across the baseline and our condition.
@@
The figure illustrates multimodal intent expression using text, sketches, and images together. User sketches provide spatial structure, while textual annotations specify narrative details. The system integrates these inputs to generate aligned image–text cards reflecting the intended scene.
A.3. Writing Topics Used in the User Study